Circular No. 8066

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002kk AND 2003ae Further to IAUC 8060, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and B. Farris, Lawrence Berkeley National Laboratory (LBNL), report the discovery of an apparent supernova (mag 19.4) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on 2002 Aug. 15.44, 15.45, and 15.46 UT. SN 2002kk is located at α = 21^{h}58^{m}58^{s}.65, δ = −0^{o}08′23′′.5 (equinox 2000.0). Additional unfiltered CCD magnitude estimates for SN 2002kk: 2002 Aug. 5, [20.5 (S/N = 3); 18, 19.2; 26, 19.0; Sept. 5, 19.5; 27, 20.1; 28, 20.2; Oct. 8, 20.8; 13, [20.5; 22, [20.5; Nov. 15, [20.5. There is possible evidence for a faint underlying object (at mag 22.0, S/N = 3.6) at the position of SN 2002kk in a stack of 22 images taken during 2001 Aug. 10–2002 July 12; otherwise, no host galaxy was specified.

L. Wang, G. Aldering, S. Perlmutter, and W. M. Wood-Vasey, LBNL, report that a spectrum (range 450–1000 nm) of SN 2002kk, obtained on 2002 Nov. 9 UT with the Keck II telescope (+ ESI spectrograph), showed it to be a type-Ia supernova about two months after maximum. The spectrum is consistent with a spectrum of SN 1991T taken at 67 days past maximum. This approximate date is consistent with the date of maximum derived from the unfiltered photometry, which gives Aug. 26 ± 2 days for the maximum (or 74 days before the spectrum was taken). Based on an overall χ^{2} minimization fit of the spectrum compared with SN 1991T, Wang et al. find a redshift of z = 0.09 for SN 2002kk.

Wood-Vasey, Aldering, and Nugent also report the discovery of an apparent supernova (mag 17.9) on unfiltered Palomar NEAT images taken on 2003 Jan. 23.38 and 23.40 UT. SN 2003ae is located at α = 9^{h}28^{m}22^{s}.58, δ = +27^{o}26′40′′.7 (equinox 2000.0), which is 0′′.3 west and 0′′.3 north of the center of the apparent host galaxy. Additional magnitudes for SN 2003ae: 2002 Dec. 14, [20.5 (S/N = 3); 2003 Jan. 10, [20.5; Jan. 24, 17.6; 28, 17.1.

Further to IAUC 8064, T. Matheson, P. Challis, and R. Kirshner report that a spectrum (range 370–750 nm) of SN 2003ae, obtained by M. Calkins on Feb. 6.29 UT with the 1.5-m telescope, shows it to be a type-Ia supernova before maximum: “Narrow emission lines superposed on the spectrum of the supernova indicate a recession velocity of 9600 km/s for the host galaxy. Using this value, the supernova expansion velocity is ∼ 13400 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 7 ± 2 days before maximum light; note that the discovery image was taken fourteen days before this spectrum. Photometry to determine the date of maximum would be most useful to constrain the rise time of type-Ia supernovae.”

2003 February 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green